# Free-form Generation Enhances Challenging Clothed Human Modeling

### v1

@@ Abstract
Achieving realistic animated human avatars requires accurate modeling of pose-dependent clothing deformations. Existing learning-based methods heavily rely on the Linear Blend Skinning (LBS) of minimally-clothed human models like SMPL to model deformation. However, these methods struggle to handle loose clothing, such as long dresses, where the canonicalization process becomes ill-defined when the clothing is far from the body, leading to disjointed and fragmented results. To overcome this limitation, we propose a novel hybrid framework to model challenging clothed humans. Our core idea is to use dedicated strategies to model different regions, depending on whether they are close to or distant from the body. Specifically, we segment the human body into three categories: unclothed, deformed, and generated. We simply replicate unclothed regions that require no deformation. For deformed regions close to the body, we leverage LBS to handle the deformation. As for the generated regions, which correspond to loose clothing areas, we introduce a novel free-form, part-aware generator to model them, as they are less affected by movements. This free-form generation paradigm brings enhanced flexibility and expressiveness to our hybrid framework, enabling it to capture the intricate geometric details of challenging loose clothing, such as skirts and dresses. Experimental results on the benchmark dataset featuring loose clothing demonstrate that our method achieves state-of-the-art performance with superior visual fidelity and realism, particularly in the most challenging cases.
@@ 1 Introduction
In this work, we revisit the task of clothed human modeling from a novel perspective. Our pivotal insight is that integrating structural priors, i.e. LBS, significantly facilitates the task, whereas relying entirely on LBS hampers flexibility. To that end, we propose a hybrid framework, leveraging the complementary advantages of LBS-based and LBS-free techniques. To achieve this, we first conduct part segmentation to categorize surface points on the human body into three types: unclothed (yellow), deformed (blue), and generated (green), as shown in Fig. 1 (a). The yellow areas represent unclothed parts (e.g. head, hands, and feet), usually not covered by garments, which need no deformation. The blue areas represent parts close to the body, where we perform LBS-based deformation [39, 40, 33, 70]. The green areas indicate loose clothing regions that deviate significantly from the body and are therefore less affected by body movements. For these regions, we introduce a free-form generator to model the dynamics. Finally, we obtain a completely clothed human by merging the three branches.
Our main contributions are summarized as follows:
  • Our hybrid framework, merging the merits of LBS deformation and free-form generator, delivers SOTA performance and enhanced visual fidelity and realism, especially in the most challenging cases.
@@ 3.1 Human Part Segmentation
Figure 2: Overview of our hybrid framework. Given an unclothed and posed body, and a specific garment type, our goal is to create a realistic clothed human. We first segment the human parts into three different regions (Sec. 3.1): unclothed parts (yellow) need no deformation, deformed parts (blue), and generated parts (green). The hybrid framework comprises two essential modules: (1) an LBS-based local deformation network (Sec. 3.2) to obtain pose-dependent deformed points $\boldsymbol{X}^{d}$ that are close to the human body, and (2) a free-form generator that focuses on generating the more loose clothing regions $\boldsymbol{X}^{g}$ (Sec. 3.3). By merging the unclothed, deformed, and generated points, we ultimately obtain the complete point cloud of a clothed human $\boldsymbol{X}$.
@@ 3.3 Free-form Generation for Loose Clothing
Structure-aware Pose Encoding. Conceptually, generating loose garments given a specific pose can be interpreted as a point cloud completion task. To better condition on human poses, we modify the off-the-shelf SpareNet [66] to be structure-aware as our generator. We first segment the human body surface into $K_{b}$ semantic parts and uniformly sample posed points $\{\boldsymbol{P}_{k}\}_{k=1}^{K_{b}}$ from these parts (with a slight abuse of notation). We also replace the original PointNet [52] with PointNet++ [53] as pose encoder $\mathcal{E}_{g}$ to extract part-wise local features $\boldsymbol{h}^{p}_{k}$ for each part, which are then fused to produce a part-based pose code $\boldsymbol{h}^{p}\in\mathbb{R}^{M_{p}}$. Note that $\boldsymbol{h}^{p}$ is shared for all vertices, while the local pose code $\boldsymbol{z}_{i}^{p}$ in Sec. 3.2 is unique for each vertex. In contrast to directly extracting global features from the overall point cloud, this structure-aware design better captures the correlation between the loose garment and the underlying skeleton.
@@ 3.4 Training
$\mathcal{L}_{cd}=\dfrac{1}{N}\sum\limits_{i=1}^{N}\min\limits_{j}\|\boldsymbol{x}_{i}-\hat{\boldsymbol{x}}_{j}\|_{2}^{2}+\dfrac{1}{M}\sum\limits_{j=1}^{M}\min\limits_{i}\|\boldsymbol{x}_{i}-\hat{\boldsymbol{x}}_{j}\|_{2}^{2},$ (10)
@@
$\mathcal{L}_{rd}=\dfrac{1}{N_{d}}\sum\limits_{i=1}^{N_{d}}\|\boldsymbol{r}_{i}^{c}\|_{2}^{2},\quad\mathcal{L}_{rg}=\sum\limits_{i=1}^{N_{d}}\dfrac{1}{N_{d}}\|\boldsymbol{z}_{i}^{g}\|_{2}^{2}+\|\boldsymbol{h}^{g}\|_{2}^{2}.$ (12)
@@
$\mathcal{L}_{c}=\dfrac{1}{N_{g}}\sum\limits_{j=1}^{N_{g}}\max\{\epsilon-d(\boldsymbol{x}^{g}_{j}),0\},$ (13)
@@ 4.2 Ablation Study
Clothing-cut Map. To assess the efficacy of the clothing-cut map, we conduct an ablation study where we remove the segmentation guidance when generation. As depicted in the left case of Fig. 7, this results in penetration artifacts, discontinuities, and missing details in the long dress. Note the dress splits around the left leg. In contrast, by introducing clothing-cut maps, the generator can model the whole dress holistically, avoiding conflicts with the deformation module and improving visual quality greatly.
@@ 5 Conclusion
We present a hybrid point-based solution for modeling challenging clothed humans, which integrates LBS deformation and free-form generation to tackle different clothing regions. To synergize the strengths of these two modules, we propose to segment the body surface into unclothed, deformed, and generated regions, yielding a clothing-cut map. Our innovative framework effectively eliminates the pant-like and inhomogeneous density artifacts in prior methods when modeling skirts and long dresses. The free-form generator provides enhanced topological flexibility and expressiveness, enabling our model to generate realistic and high-quality wrinkle details. We assess our model with varying skirt lengths, tightness, and styles, and the experimental results demonstrate the superior representational power of the proposed framework. We believe that this novel hybrid modeling opens up new possibilities in this domain. Additionally, our point-based hybrid modeling can be integrated with recent advancements in 3DGS [25] to enhance texture rendering, which we plan to explore in future work.
@@ A.1 Model Architecture
The structure-aware pose encoder $\mathcal{E}_{g}$ for extracting pose feature embedding for the free-form generation module possesses a similar architecture with $\mathcal{E}_{d}$. Given our focus on modeling skirts and long dresses, we selectively sample posed points from $\boldsymbol{K}_{b}=4$ local parts situated on the legs, including the left upper leg, left lower leg, right upper leg, and right lower leg. Specifically, we uniformly sample $2048$ points from each part, which are then inputted into $\mathcal{E}_{g}$ to derive part-aware local features. A final global max-pooling layer is prepended to extract the global pose features.
@@ A.3 Training
We train our network for $1000$ epochs on the ReSynth [39] dataset, using the Adam [27] optimizer with a batch size of $8$ and a learning rate of $3.0\times{10}^{-4}$. The loss weights are set to $\lambda_{p}=1\times{10}^{4}$, $\lambda_{n}=1.0$, $\lambda_{rd}=2\times 10^{3}$, $\lambda_{rg}=1$ and $\lambda_{col}=2\times 10^{-2}$ to balance loss terms. Following previous works [39, 70], we only activate the normal loss from the $400^{th}$ epoch. The training procedure takes about 20 hours on a single RTX 3090 GPU.
Given limited 3D training data, we enhance the robustness of our free-form generator to out-of-distribution poses by balancing the pose distribution. Specifically, we apply random horizontal flips along the $x$-axis, leveraging the symmetry of the human body.
@@ B.1 Discussions on the Evaluation Metric
Therefore, building on the limitations discussed, we employ a generation-based metric, FID, which compares distributions and relaxes the strict one-to-one mapping constraint, alongside the reconstruction-based MSE loss to assess our model’s quality holistically. This approach better aligns with the model’s objective than traditional reconstruction metrics.
@@ B.3 More Qualitative Results
In this section, we present additional visualization comparisons that extend the results discussed in the main paper. Fig. B6 illustrates the details of the generated loose clothing, which are highlighted within the red box. Furthermore, we showcase three testing examples for each of the five subjects from the ReSynth [39] dataset, as illustrated in Figs. B14, B15, B16, B17 and B18. We recommend zooming in to observe finer details, particularly the wrinkles in skirts and dresses. Please refer to the visualization demo in our supplementary materials, which includes sequences of testing data to better demonstrate the high-quality performance of our method.
As discussed in the main paper, the advantages of our hybrid modeling approach become particularly evident with loose skirts or dresses, as illustrated by the examples in Figs. B14 and B15. For tighter skirts, LBS-based models like FITE [33] already perform well since the clothing adheres closely to the body. As shown in Figs. B17 and B18, FITE [33] generates nearly perfect outputs that closely resemble the ground truth, and our model produces results comparable to those of FITE [33]. However, it is noteworthy that FITE [33] still fails to fully eliminate redundant points on the open surface of tighter skirts (see Fig. B7). This represents a limitation compared to our method.
@@
B.4 Multi-Subject Experiments
In this study, we explore the potential of hybrid modeling for loose clothing and significantly improve the performance under a single-subject setting, compared with previous works [40, 33, 70]. Nevertheless, our hybrid paradigm can also be naturally extended to modeling multiple garments, conditioned on various global garment codes.
Surprisingly, our unified, multi-subject model demonstrates promising performance in modeling various types of skirts and long dresses, confirming the expressive power of our free-form generator. To explore the learned latent space of garment codes, we perform interpolation experiments focusing on two crucial attributes: length and tightness. As shown in Fig. B8, our model allows effective control over garment length through manipulation of the garment code. Furthermore, when varying the length, the generated skirts smoothly transition from tight to loose. In summary, our model successfully disentangles pose-related effects from garment-specific features, providing controllable and realistic generation results.
@@ B.5 Fitting Non-skirt Clothing
Figure B9: Our hybrid model can also handle non-skirt clothing such as suits. As shown on the left-hand side, the free-form generation module can model the loose regions such as collars.
@@ B.6 More Ablation Studies
The visualization result (a) shows the inadequacy of setting $K=2$ for capturing the intricate details of the long dress, resulting in over-smoothed outcomes. In addition, empirical findings indicate that the setting $K=8$ is sufficient to generate high-quality details, producing superior visual results. When $K$ surpasses $8$ or is set to $4$, the generated surface exhibits increased noise, leading to a loss of clarity in fine-grained details. Notably, features such as clothing wrinkles become less distinct and sharply defined. This observation suggests that modeling the ostensibly complex long dress may be less daunting than anticipated. Furthermore, it verifies the remarkable expressiveness of our hybrid framework. Unless otherwise stated, $K=8$ is selected for our experiments.
@@
Figure B11: Visualization of the generated $K=8$ patches which comprise the loose dress and skirts, marked in different colors.
Figure B12: More ablation study results of employing clothing-cut map. Split-up artifacts between two modules can manifest under various poses, especially when the leg approaches the dress surface, resulting in disjointed, noisy areas and torn appearances. The use of the clothing-cut map notably mitigates this issue.

### v2

# FreeCloth: Free-form Generation Enhances Challenging Clothed Human Modeling

@@ Abstract
Achieving realistic animated human avatars requires accurate modeling of pose-dependent clothing deformations. Existing learning-based methods heavily rely on the Linear Blend Skinning (LBS) of minimally-clothed human models like SMPL to model deformation. However, they struggle to handle loose clothing, such as long dresses, where the canonicalization process becomes ill-defined when the clothing is far from the body, leading to disjointed and fragmented results. To overcome this limitation, we propose FreeCloth, a novel hybrid framework to model challenging clothed humans. Our core idea is to use dedicated strategies to model different regions, depending on whether they are close to or distant from the body. Specifically, we segment the human body into three categories: unclothed, deformed, and generated. We simply replicate unclothed regions that require no deformation. For deformed regions close to the body, we leverage LBS to handle the deformation. As for the generated regions, which correspond to loose clothing areas, we introduce a novel free-form, part-aware generator to model them, as they are less affected by movements. This free-form generation paradigm brings enhanced flexibility and expressiveness to our hybrid framework, enabling it to capture the intricate geometric details of challenging loose clothing, such as skirts and dresses. Experimental results on the benchmark dataset featuring loose clothing demonstrate that FreeCloth achieves state-of-the-art performance with superior visual fidelity and realism, particularly in the most challenging cases.
@@ 1 Introduction
In this work, we revisit the task of clothed human modeling from a novel perspective. Our pivotal insight is that integrating structural priors, i.e. LBS, significantly facilitates the task, whereas relying entirely on LBS hampers flexibility. To that end, we propose FreeCloth, a hybrid framework, leveraging the complementary advantages of LBS-based and LBS-free techniques. We first conduct part segmentation to categorize surface points on the human body into three types: unclothed (yellow), deformed (blue), and generated (green), as shown in Fig. 1 (a). The yellow areas represent unclothed parts (e.g. head, hands, and feet), usually not covered by garments, which need no deformation. The blue areas represent parts close to the body, where we perform LBS-based deformation [39, 40, 33, 70]. The green areas indicate loose clothing regions that deviate significantly from the body and are therefore less affected by body movements. For these regions, we introduce a free-form generator to model the dynamics. Finally, we obtain a completely clothed human by merging the three branches.
@@
Our main contributions are summarized as follows:
  • Our hybrid framework, FreeCloth, merging the merits of LBS deformation and free-form generator, delivers SOTA performance and enhanced visual fidelity and realism, especially in the most challenging cases.
@@ 3 Method
Figure 2: Overview of our hybrid framework FreeCloth. Given an unclothed and posed body, and a specific garment type, our goal is to create a realistic clothed human. We first segment the human parts into three different regions (Sec. 3.1): unclothed parts (yellow) need no deformation, deformed parts (blue), and generated parts (green). The hybrid framework comprises two essential modules: (1) an LBS-based local deformation network (Sec. 3.2) to obtain pose-dependent deformed points $\boldsymbol{X}^{d}$ that are close to the human body, and (2) a free-form generator that focuses on generating the more loose clothing regions $\boldsymbol{X}^{g}$ (Sec. 3.3). By merging the unclothed, deformed, and generated points, we ultimately obtain the complete point cloud of a clothed human $\boldsymbol{X}$.
@@ 3.3 Free-form Generation for Loose Clothing
Structure-aware Pose Encoding. Conceptually, generating loose garments given a specific pose can be interpreted as a task of point cloud completion. To better condition on human poses, we modify the off-the-shelf SpareNet [66] to be structure-aware as our generator. We first segment the human body surface into $K_{b}$ semantic parts and uniformly sample posed points $\{\boldsymbol{P}_{k}\}_{k=1}^{K_{b}}$ from these parts (with a slight abuse of notation). We also replace the original PointNet [52] with PointNet++ [53] as pose encoder $\mathcal{E}_{g}$ to extract part-wise local features $\boldsymbol{h}^{p}_{k}$ for each part, which are then fused to produce a part-based pose code $\boldsymbol{h}^{p}\in\mathbb{R}^{M_{p}}$. Note that $\boldsymbol{h}^{p}$ is shared for all vertices, while the local pose code $\boldsymbol{z}_{i}^{p}$ in Sec. 3.2 is unique for each vertex. In contrast to directly extracting global features from the overall point cloud, this structure-aware design better captures the correlation between the loose garment and the underlying skeleton.
@@ 3.4 Training
$\small\mathcal{L}_{cd}=\dfrac{1}{N}\sum\limits_{i=1}^{N}\min\limits_{j}\|\boldsymbol{x}_{i}-\hat{\boldsymbol{x}}_{j}\|_{2}^{2}+\dfrac{1}{M}\sum\limits_{j=1}^{M}\min\limits_{i}\|\boldsymbol{x}_{i}-\hat{\boldsymbol{x}}_{j}\|_{2}^{2},$ (10)
@@
$\small\mathcal{L}_{rd}=\dfrac{1}{N_{d}}\sum\limits_{i=1}^{N_{d}}\|\boldsymbol{r}_{i}^{c}\|_{2}^{2},\quad\mathcal{L}_{rg}=\sum\limits_{i=1}^{N_{d}}\dfrac{1}{N_{d}}\|\boldsymbol{z}_{i}^{g}\|_{2}^{2}+\|\boldsymbol{h}^{g}\|_{2}^{2}.$ (12)
@@
$\small\mathcal{L}_{c}=\dfrac{1}{N_{g}}\sum\limits_{j=1}^{N_{g}}\max\{\epsilon-d(\boldsymbol{x}^{g}_{j}),0\},$ (13)
@@ 4.2 Ablation Study
Clothing-cut Map. To assess the efficacy of the clothing-cut map, we conduct an ablation study by removing segmentation guidance and reverting to the default UV map. In this setting, body points in loose regions are also deformed, leading to the blending of loose clothing points from two branches. As depicted in Fig. 7, this results in penetration artifacts, discontinuities, and missing details in the long dress. Note the dress splits around the left leg. In contrast, by introducing clothing-cut maps, the generator can model the whole dress holistically, avoiding conflicts with the deformation module and improving visual quality greatly.
@@ 5 Conclusion
We present FreeCloth, a hybrid point-based solution for modeling challenging clothed humans, which integrates LBS deformation and free-form generation to tackle different clothing regions. To synergize the strengths of these two modules, we propose to segment the body surface into unclothed, deformed, and generated regions, yielding a clothing-cut map. Our innovative framework effectively eliminates the pant-like and inhomogeneous density artifacts in prior methods when modeling skirts and long dresses. The free-form generator provides enhanced topological flexibility and expressiveness, enabling our model to generate realistic and high-quality wrinkle details. We assess our model with varying skirt lengths, tightness, and styles, and the experimental results demonstrate the superior representational power of the proposed framework. We believe that this novel hybrid modeling opens up new possibilities in this domain. Additionally, our point-based hybrid modeling can be integrated with recent advancements in 3DGS [25] to enhance texture rendering, which we plan to explore in future work.
Acknowledgment
This work was supported by 2022ZD0114900 and NSFC-6247070125.
@@ A.1 Model Architecture
The structure-aware pose encoder $\mathcal{E}_{g}$ for extracting pose feature embedding for the free-form generation module possesses a similar architecture with $\mathcal{E}_{d}$. Given our focus on modeling skirts and long dresses, we selectively sample posed points from $\boldsymbol{K}_{b}=4$ local parts situated on the legs, including the left upper leg, left lower leg, right upper leg, and right lower leg. Although the short skirt doesn’t directly cover the lower legs, their pose still indirectly affects the skirt’s movement. Specifically, we uniformly sample $2048$ points from each part, which are then inputted into $\mathcal{E}_{g}$ to derive part-aware local features. A final global max-pooling layer is prepended to extract the global pose features.
@@ A.3 Training
We train our network for $1000$ epochs on the ReSynth [39] dataset, using the Adam [27] optimizer with a batch size of $8$ and a learning rate of $3.0\times{10}^{-4}$. The loss weights are set to $\lambda_{p}=1\times{10}^{4}$, $\lambda_{n}=1.0$, $\lambda_{rd}=2\times 10^{3}$, $\lambda_{rg}=1$ and $\lambda_{col}=2\times 10^{-2}$ to balance loss terms. Following previous works [39, 70], we only activate the normal loss from the $400^{th}$ epoch. The training procedure takes about 20 hours on a single RTX 3090 GPU. Given limited 3D training data, we enhance the robustness of our free-form generator to out-of-distribution poses by balancing the pose distribution. Specifically, we apply random horizontal flips along the $x$-axis, leveraging the symmetry of the human body.
@@ B.1 Discussions on the Evaluation Metric
Building on the limitations discussed, we employ a generation-based metric, FID, which compares distributions and relaxes the strict one-to-one mapping constraint, alongside the reconstruction-based MSE loss to assess our model’s quality holistically. This evaluation approach better aligns with the model’s objective.
@@ B.3 More Qualitative Results
In this section, we present additional visualization comparisons that extend the results discussed in the main paper. Fig. B6 illustrates the details of the generated loose clothing, which are highlighted within the red box. Furthermore, we showcase three testing examples for each of the five subjects from the ReSynth [39] dataset, as illustrated in Figs. B15, B16, B17, B18 and B19. We recommend zooming in to observe finer details, particularly the wrinkles in skirts and dresses. Please refer to the visualization demo in our supplementary materials, which includes sequences of testing data to better demonstrate the high-quality performance of our method.
As discussed in the main paper, the advantages of our hybrid modeling approach become particularly evident with loose skirts or dresses, as illustrated by the examples in Figs. B15 and B16. For tighter skirts, LBS-based models like FITE [33] already perform well since the clothing adheres closely to the body. As shown in Figs. B18 and B19, FITE [33] generates nearly perfect outputs that closely resemble the ground truth, and our model produces results comparable to those of FITE [33]. However, it is noteworthy that FITE [33] still fails to fully eliminate redundant points on the open surface of tighter skirts (see Fig. B7).
@@
B.4 Multi-Subject Experiments
In this study, we explore the potential of hybrid modeling for loose clothing and significantly improve the performance under a single-subject setting. Nevertheless, our hybrid paradigm can be naturally extended to modeling multiple garments, conditioned on various global garment codes. Experimental results show that our unified, multi-subject model demonstrates promising performance in modeling various types of skirts and long dresses, confirming the expressive power of our free-form generator.
@@ B.4 Multi-Subject Experiments
To explore the learned latent space of garment codes, we perform interpolation experiments focusing on two crucial attributes: length and tightness. As shown in Fig. B8, our model allows effective control over garment length through manipulation of the garment code. Furthermore, when varying the tightness, the generated skirts smoothly transition from tight to loose. In summary, our model successfully disentangles pose-related effects from garment-specific features, providing controllable and realistic generation results.
@@ B.5 Fitting Non-skirt Clothing
Table B3: Ablation study of the free-form generation module on the ReSynth [39] dataset. In the setting Ours, the free-form generator is removed, relying solely on body point deformation to model loose clothing.
Ours 56.23 2.73 63.12 5.72 52.10 2.06 59.29 2.41 51.68 1.84 54.96 1.62
Ours 37.75 2.61 42.41 5.24 27.95 1.92 37.43 2.35 39.63 1.89 41.24 1.68
Figure B9: Our hybrid model can also handle non-skirt clothing such as suits. As shown on the left-hand side, the free-form generation module can model loose regions such as collars.
@@ B.6 More Ablation Studies
Effects of the Hybrid Paradigm.
To evaluate the efficacy of the proposed free-form generator, we quantitatively assess the deformation-only variant using the ReSynth [39] dataset. As shown in Tab. B3, this variant achieves an average FID of 56.23 and MSE of 2.74, comparable to SkiRT [40]. In contrast, our full model substantially improves these metrics, demonstrating the effectiveness of the free-form generation module in capturing the dynamics of loose clothing. Additionally, we examine a generation-only variant that discards LBS-based deformation and synthesizes full-body clothing points from a global pose feature, as illustrated in Fig. B10. The resulting noisy surfaces in articulated regions and overly coarse details further emphasize the necessity of the hybrid paradigm.
Figure B10: Ablation study of the full-body free-form generation. (a) Completely discarding LBS deformation results in a significant performance drop when compared to (b) our full model.
@@
Visualization (a) shows that $K=2$ fails to capture the intricate details of the long dress, resulting in over-smoothed outcomes. In addition, empirical findings indicate that the setting $K=8$ is sufficient to generate high-quality details, producing superior visual results. When $K$ surpasses $8$ or is set to $4$, the generated surface exhibits increased noise, leading to a loss of clarity in fine-grained details. Notably, features such as clothing wrinkles become less distinct and sharply defined. This observation suggests that modeling the ostensibly complex long dress may be less daunting than anticipated. Furthermore, it verifies the remarkable expressiveness of our hybrid framework. Unless otherwise stated, $K=8$ is selected for our experiments.
@@
Figure B12: Visualization of the generated $K=8$ patches which comprise the loose dress and skirts. We show the results on the three subjects.
Figure B13: More ablation study results of employing clothing-cut map. Split-up artifacts between two modules can result in disjointed, noisy areas and torn appearances. The use of the clothing-cut map notably mitigates this issue.
@@
B.7 Efficiency Analysis
We evaluate the inference speed of our approach and other SOTA methods on an RTX 3090 GPU, with a batch size of 1. Additionally, we report the FLOPs and parameter counts to quantify computational resource requirements. As shown in Tab. B4, our model has the smallest number of parameters and achieves real-time inference speed at $64.1$ FPS. Notably, our model significantly outperforms SOTAs without introducing extra computational overhead.
Table B4: Efficiency analysis of our method with other works.
Method FID $\downarrow$ FPS $\uparrow$ FLOPS (G) Params. (M)
POP [39] 57.87 69.9 128.81 11.33
SkiRT [40] 53.32 79.9 77.12 11.13
FITE [33] 39.02 31.5 68.87 11.02
Ours 37.75 64.1 78.82 10.83
B.8 Limitations and Failure Cases
Despite employing pose augmentation, our model remains susceptible to failure when confronted with extremely challenging poses, resulting in clothing penetration artifacts, as depicted in Fig. B14 (a). This issue is particularly noticeable when the skirt becomes tighter. Training our free-form generator on a larger dataset could enhance its robustness to out-of-distribution poses and reduce such artifacts. Additionally, while our pipeline employs different strategies to handle deformed and generated areas, we cannot guarantee the perfect blending of point clouds from two modules. As illustrated in Fig. B14 (b), “seams” at the boundary regions are occasionally observed.